# Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v1

@@ Abstract
Label smoothing (LS) is a popular regularisation method for training deep neural network classifiers due to its effectiveness in improving test accuracy and its simplicity in implementation. “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting. In this work, we reveal that LS negatively affects selective classification (SC) – where the aim is to reject misclassifications using a model’s predictive uncertainty. We first demonstrate empirically across a range of tasks and architectures that LS leads to a consistent degradation in SC. We then explain this by analysing logit-level gradients, showing that LS exacerbates overconfidence and underconfidence by regularising the max logit more when the probability of error is low, and less when the probability of error is high. This elucidates previously reported experimental results where strong classifiers underperform in SC. We then demonstrate the empirical effectiveness of logit normalisation for recovering lost SC performance caused by LS. Furthermore, based on our gradient analysis, we explain why such normalisation is effective. We will release our code shortly.
Keywords:
Uncertainty Estimation Selective Classification Label Smoothing
@@ 1 Introduction
Label smoothing (LS) [70] is a common regularisation technique used to improve classification accuracy in supervised deep learning. When training with cross-entropy minimisation, one-hot labels are linearly combined with a uniform distribution over classes, redistributing the probability mass and “smoothing” the “hard” targets. This discourages overfitting on the training labels. Due to its simplicity and empirical effectiveness, label smoothing features in many recent training recipes [75, 71, 28, 74, 52, 51, 72, 73, 53], being particularly popular on the ImageNet-1k [63] image classification benchmark.
@@
Within the domain of uncertainty estimation, label smoothing is well explored in the context of model calibration [59, 10, 58, 47], where the aim is to align a model’s output probabilities with its empirical accuracy. The pairing is intuitive as label smoothing encourages models to be less confident, and models are typically more confident than they are accurate [25]. On the other hand, there is a distinct lack of research investigating the effects of label smoothing in the context of Selective Classification. Selective Classification (SC) [30, 22, 81, 36] is a problem setting where, in addition to the primary classification task, a binary rejection decision is made based on the predictive uncertainty estimated by the model, i.e. reject if uncertain. The aim is to reduce the number of errors served by the classifier by pre-emptively rejecting potential misclassifications. It is well motivated by applications where safety and reliability are important due to the high cost of failure, such as autonomous driving [37] and medical diagnosis [1, 42].
[…]
Recent large-scale empirical evaluations of pre-trained models [20, 4] have shown that many strong classifiers have surprisingly poor SC ability. That is to say, in spite of having very low error rates, they are poor at rejecting their own misclassifications. Unfortunately, by evaluating pre-trained models from open-source repositories, the authors find it difficult to disentangle any potential causes of this behaviour, as model training recipes vary wildly from checkpoint to checkpoint. In this work, we reveal that label smoothing is one such potential cause and present the following key contributions:
@@
    We show empirically, across a range of network architectures and vision tasks, that training with LS consistently leads to degraded SC performance, even if it may improve accuracy (see Fig. 1). As LS can be found in the training recipes of many of the models evaluated in [20, 4], this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
@@
    We are able to explain the above behaviour by analysing the logit-level gradients of the LS loss. We show that the amount LS regularises the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing as $P_{\text{error}}$ decreases. This exacerbates overconfidence and underconfidence, degrading SC.
@@
    Finally, we demonstrate that logit normalisation proposed by Cattelan and Silva [4] is highly empirically effective in combating specifically the degradation caused by LS. Moreover, we are able to explain the effectiveness of logit normalisation through the lens of our gradient-based analysis, showing that normalisation compensates for the imbalanced regularisation by increasing uncertainty as the max logit increases. [^1] Although they demonstrate its empirical effectiveness, Cattelan and Silva [4] do not make a connection to label smoothing, and provide limited insight as to why their approaches are effective.
@@ 2 Preliminaries
$P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})=\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=\exp v_{k}(\boldsymbol{x})/\sum_{i=1}^{K}\exp v_{i}(\boldsymbol{x})~{},\quad\boldsymbol{v}=\boldsymbol{W}\boldsymbol{z}+\boldsymbol{b}~{},$ (1)
where $\boldsymbol{v}\in\mathbb{R}^{K}$ are the logits output by the final layer with weight matrix $\boldsymbol{W}\in\mathbb{R}^{K\times L}$, bias $\boldsymbol{b}\in\mathbb{R}^{K}$, and pre-logit features $\boldsymbol{z}\in\mathbb{R}^{L}$ as inputs. The classifier is trained by minimising the cross entropy (CE) loss on a finite dataset $\mathcal{D}_{\text{tr}}=\{x^{(n)},y^{(n)}\}_{n=1}^{N}$ drawn from distribution $p_{\text{data}}(\boldsymbol{x},y)$, such that it approximately learns the true conditional $P_{\text{data}}(y|\boldsymbol{x})$
$\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ $=-\frac{1}{N}\sum_{n}\sum_{k}\delta_{y^{(n)}\omega_{k}}\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (2)
$\approx-\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\biggl{[}\sum_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})\log P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})\biggr{]}$ (3)
@@
where $\delta$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence. We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e. $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$. Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$,
$\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})~{}.$ (5)
We also define the probability of the classifier making an error as,
$P_{\text{error}}=1-\bar{\pi}_{\hat{y}}~{},$ (6)
where, in a slight abuse of notation $\bar{\pi}_{\hat{y}}$ is the true probability of the predicted class.
@@ 2.1 Label Smoothing (LS)
Figure 2: Left: illustration of how label smoothing (LS) alters a training label. LS reduces data supervision and adds regularisation. Right: illustration of selective classification (SC). Uncertain samples ($U>\tau$) are rejected, with the aim of reducing the number of errors  served.
Label smoothing is most commonly understood (and implemented) as mixing the original one-hot labels (represented by $\delta$ in Eq. 2) with a uniform categorical distribution $\boldsymbol{u}=\boldsymbol{1}/K$ using hyperparameter $\alpha\in[0,1]$ (see Fig. 2). The label smoothing loss is thus
$\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ $=-\frac{1}{N}\sum_{n}\sum_{k}\left[(1-\alpha)\delta_{y^{(n)}\omega_{k}}+\alpha\frac{1}{K}\right]\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (7)
$=(1-\alpha)\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})+\alpha\frac{1}{N}\sum_{n}\left[\text{KL}\left[\boldsymbol{u}||\boldsymbol{\pi}(\boldsymbol{x}^{(n)};\boldsymbol{\theta})\right]\right]+\text{const.}~{}$. (8)
@@
where we see that it can also be viewed as reduced CE supervision from the data combined with a regularisation term encouraging the softmax $\boldsymbol{\pi}$ to be uniform and preventing it from overfitting to the training data (Fig. 2). Eq. 9 also shows that LS can be seen as learning to predict a “softened” version of the true conditional $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$.
@@ 2.2 Selective Classification (SC)
A simple downstream task for estimates of predictive uncertainty is to reject (or detect) predictions that may incur a high cost [36, 81], using a binary rejection function,
@@
where $U(\boldsymbol{x})$ is a scalar measure of predictive uncertainty extracted from the prediction model and $\tau$ is the operating threshold. Intuitively, we reject if the model is uncertain. Notably, we are only concerned with the relative rankings of $U$s rather than the absolute values. This differs from model calibration [25, 36], where the value of uncertainty predicted should match the marginal error rate over the data distribution for that value.
In the case where the prediction task is classification and we wish to reject potential misclassifications (), we can use a selective classifier [14, 9] $(f,g)$, which is simply the combination of a classifier $f$ (Eq. 5) and the aforementioned binary rejection function $g$ (Eq. 10). Fig. 2 contains an illustration. Given the 0/1 classification error,
@@
we define the selective risk [14, 22] as
@@
$U(\boldsymbol{x})=-\text{MSP}(\boldsymbol{x})=-\pi_{\text{max}}=-\max_{k}\pi_{k}(\boldsymbol{x};\theta)=-P(\hat{y}|\boldsymbol{x};\boldsymbol{\theta})~{},$ (13)
i.e. the (negative of the) Maximum Softmax Probability (MSP) [30] is the model’s estimate of the probability of its prediction $\hat{y}$. We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature [64, 81, 83, 36, 4, 15, 38, 22, 30].
2.2.1 Overconfidence.
@@
Since we are only concerned about the rank ordering of uncertainties $U$, we loosely refer to overconfidence as when a model’s estimate of uncertainty is (relatively) low when $P_{\text{error}}$ high, as intuitively we would like a model to be uncertain when it is likely to be wrong. Underconfidence is then the inverse. Both will result in worse SC performance as overconfidence will cause errors to be accepted whilst underconfidence will lead to correct predictions being rejected. This is different to the more specific definition used in the problem scenario of model calibration [25].
2.3 Aleatoric and Epistemic Uncertainty
Figure 3: Left: different types of uncertainty for a model trained on ImageNet [63]. Aleatoric uncertainty is high as a different label $y$ could be sampled, while epistemic uncertainty is high as the image is from an unfamiliar domain, here ImageNet-Sketch [76]. Right: how errors arise from different uncertainties. Aleatoric uncertainty leads to errors as there are many probable labels. Epistemic uncertainty leads to errors as the model is unable to predict the true distribution $\bar{\boldsymbol{\pi}}$.
It is useful to separate uncertainty into two different types [39, 19, 35]. Aleatoric (or data) uncertainty is irreducible uncertainty in the data distribution $P_{\text{data}}(y|\boldsymbol{x})$, e.g. when there are multiple valid labels. Epistemic (or knowledge) uncertainty arises from a model’s lack of knowledge about data. A model should have greater epistemic uncertainty on the tails or low-density regions of $p_{\text{data}}(\boldsymbol{x})$, as these regions of the data distribution will be weighted less in the training loss (Eq. 4). There will also be epistemic uncertainty for different images not in training data, or images from a different domain. These different uncertainties are illustrated visually in Fig. 3 (left). While errors caused by aleatoric uncertainty are unavoidable, those due to epistemic uncertainty arise from the model’s inability to accurately predict the true conditional distribution $\bar{\boldsymbol{\pi}}$. These can be avoided with additional supervision. Both types of errors are illustrated in Fig. 3 (right). Epistemic uncertainty also intuitively results in both overconfidence and underconfidence. As the neural network models the data less accurately, its uncertainty estimates also more poorly reflect $P_{\text{error}}$ (both cases in Fig. 3 (right) are overconfident).
2.4 Experimental Details
We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Appendix 0.A. image classification on ImageNet-1k [63] and semantic segmentation (pixel-level classification) on Cityscapes [12]. For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation. For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images. To estimate the risk, we extract the model output prior to the final interpolation and subsample 5000 labelled pixels per image at random. We evaluate on the same pixels between models. All experiments are performed in Pytorch [61] and we will release our code shortly.
2.4.1 Models and Training

### v2

# Towards Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

@@ Abstract
Label smoothing (LS) is a popular regularisation method for training neural networks as it is effective in improving test accuracy and is simple to implement. “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting. Prior work has suggested that in some cases LS can degrade selective classification (SC) – where the aim is to reject misclassifications using a model’s uncertainty. In this work, we first demonstrate empirically across an extended range of large-scale tasks and architectures that LS consistently degrades SC. We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by regularising the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong. This elucidates previously reported experimental results where strong classifiers underperform in SC. We then demonstrate the empirical effectiveness of post-hoc logit normalisation for recovering lost SC performance caused by LS. Furthermore, linking back to our gradient analysis, we again provide an explanation for why such normalisation is effective.
@@ 1 Introduction
Label smoothing (LS) (Szegedy et al., 2016) is a common regularisation technique used to improve classification accuracy in deep learning. The one-hot labels used for cross entropy (CE) are linearly combined with a uniform distribution over classes, redistributing the probability mass and “smoothing” the “hard” targets, discouraging overfitting. Due to its simplicity and empirical effectiveness, label smoothing features in many recent training recipes (Vaswani et al., 2017; Tan et al., 2019; He et al., 2019; Touvron et al., 2021; Liu et al., 2021; 2022b; 2022c; Tan & Le, 2019; 2021), being particularly popular on the ImageNet-1k (Russakovsky et al., 2015) image classification benchmark.
@@
Within the domain of uncertainty estimation, LS is well explored in the context of model calibration (Müller et al., 2019; Chun et al., 2020; Mukhoti et al., 2020; Liu et al., 2022a), where the aim is to align a model’s output probabilities with its empirical accuracy. The pairing is intuitive as LS encourages models to output lower probabilities, and models are typically more confident than they are accurate (Guo et al., 2017). On the other hand, there is very little research investigating the effects of LS in the context of Selective Classification. Selective Classification (SC) (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017; Xia & Bouganis, 2022b; Jaeger et al., 2023) is a problem setting where, in addition to the primary classification task, a binary rejection decision is made based on the uncertainty estimated by the model, i.e. reject/abstain if uncertain. The aim is to reduce the number of failures served by the classifier by pre-emptively rejecting potential misclassifications. It is well motivated by applications where safety and reliability are important due to the high cost of failure. For example, when uncertain, an autonomous driving system may require driver assistance  (Kendall & Gal, 2017), a medical diagnosis system may defer to a doctor (Beam & Kompa, 2021; Kurz et al., 2022), or a visual aid system for the visually impaired may abstain from answering a query (Whitehead et al., 2022).
ResNet-50 ImageNet MSP CE LS $\alpha=0.1$ LS $\alpha=0.2$ LS $\alpha=0.3$ LS $\alpha=0.2$ – logit norm
%coverage@risk=1$\uparrow$ 17.69 0.2 0.53 0.49 21.9
Figure 1: Top: LS causes overconfidence for semantic segmentation. The LS-trained model predicts much lower (ranked) uncertainty on incorrect ✗ segmentations than CE. In particular, for the erroneous region on the left where the model has predicted parts of the “sidewalk” as “road”, the LS model is highly overconfident. This could have dire consequences in a safety-critical application such as autonomous driving. Bottom: LS leads to close to 0% of samples being accepted (coverage) when a strict tolerance of 1% error on accepted samples (risk) is imposed on ImageNet. Deployment-time logit normalisation effectively negates the degradation caused by LS.
Recently Zhu et al. (2022; 2024) empirically observe that LS can lead to worse SC for convolutional neural networks (CNNs) performing image classification, yet it is not clear why this occurs. Concurrently, large-scale empirical evaluations of open-source pre-trained models have shown that many strong classifiers have surprisingly poor SC ability (Galil et al., 2023; Cattelan & Silva, 2024). Cattelan & Silva (2024) additionally present deployment-time logit normalisation as a sometimes-effective approach to improving SC, but the reason for its effectiveness remains unclear. In this work, we aim to empirically validate and analytically demystify the effect of LS on SC, tying together and filling in the gaps in knowledge of previous work with the following key contributions:
@@
    We show empirically, across a range of large-scale architectures (CNN,ViT) and tasks (image classification, semantic segmentation), that training with LS consistently leads to degraded SC performance (see Fig. 1), even if it may improve accuracy. Moreover, we show that the degradation worsens with stronger LS. As LS can be found in the training recipes of many of the models evaluated in (Galil et al., 2023; Cattelan & Silva, 2024), this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
@@
    We address a gap in the understanding of LS by providing an explanation of this behaviour through analysing the logit-level gradients of the LS loss. We show that the amount LS regularises/suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing(decreasing) the more likely a prediction is correct(wrong). This leads to relatively higher uncertainty on correct predictions and lower uncertainty on misclassifications, degrading the ranking of uncertainties and hurting SC compared to vanilla CE.
@@
    We show that post-hoc logit normalisation (Cattelan & Silva, 2024) is effective in negating the degradation from LS (Fig. 1). Moreover, we elucidate this effectiveness through the lens of our gradient-based analysis. Linking back to the imbalanced logit regularisation of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
@@ 2 Preliminaries
$P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})=\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=\exp v_{k}(\boldsymbol{x})/\sum\nolimits_{i=1}^{K}\exp v_{i}(\boldsymbol{x})~{},\quad\boldsymbol{v}=\boldsymbol{W}\boldsymbol{z}+\boldsymbol{b}~{},$ (1)
where $\boldsymbol{v}\in\mathbb{R}^{K}$ are the logits output by the final layer with weight matrix $\boldsymbol{W}\in\mathbb{R}^{K\times L}$, bias $\boldsymbol{b}\in\mathbb{R}^{K}$, and pre-logit features $\boldsymbol{z}\in\mathbb{R}^{L}$ as inputs. The neural network is trained by minimising the cross entropy (CE) loss on a finite dataset $\mathcal{D}_{\text{tr}}=\{x^{(n)},y^{(n)}\}_{n=1}^{N}$ drawn from distribution $p_{\text{data}}(\boldsymbol{x},y)$, such that it approximately learns the true conditional $P_{\text{data}}(y|\boldsymbol{x})$,
$\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ $=-\frac{1}{N}\sum\nolimits_{n}\sum\nolimits_{k}\delta_{y^{(n)}\omega_{k}}\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (2)
$\approx-\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\biggl{[}\sum\nolimits_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})\log P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})\biggr{]}$ (3)
@@
where $\delta$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence. We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e. $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$. Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$ during deployment,
$\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}\nolimits_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})~{}.$ (5)
We also define the probability of the classifier making an error on a given sample as $P_{\text{error}}=1-\bar{\pi}_{\hat{y}}~{}$, where in a slight abuse of notation $\bar{\pi}_{\hat{y}}$ is the true probability of the predicted class $P_{\text{data}}(\hat{y}|\boldsymbol{x})$.
@@ 2.1 Label smoothing (LS)
Figure 2: Left: illustration of how label smoothing (LS) alters a training label. LS reduces data supervision and adds regularisation, potentially improving generalisation by reducing overfitting. Right: illustration of selective classification (SC). Uncertain samples ($U>\tau$) are rejected/detected, to reduce the number of errors  served by the system. Rejected samples can be discarded or processed separately (e.g. deferred to a human expert). We wish to better separate/rank  vs  via $U$.
Label smoothing involves mixing the original one-hot labels ($\delta$ in Eq. 2) with a uniform categorical distribution $\boldsymbol{u}=1/K\cdot\boldsymbol{1}$ using hyperparameter $\alpha\in[0,1]$ (see Fig. 2). The LS loss is thus
$\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ $=-\frac{1}{N}\sum\nolimits_{n}\sum\nolimits_{k}\left[(1-\alpha)\delta_{y^{(n)}\omega_{k}}+\alpha\frac{1}{K}\right]\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (6)
$=(1-\alpha)\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})+\alpha\frac{1}{N}\sum\nolimits_{n}\left[\text{KL}\left[\boldsymbol{u}||\boldsymbol{\pi}(\boldsymbol{x}^{(n)};\boldsymbol{\theta})\right]\right]+\text{const.}~{}$. (7)
@@
where we see that it can also be viewed as reduced CE supervision from the data combined with a regularisation term encouraging the softmax $\boldsymbol{\pi}$ to be uniform and preventing it from overfitting to the training data (Fig. 2). Eq. 8 also shows that LS can be seen as learning to predict a “softened” version of the true conditional $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$, encouraging a model to be less confident on all samples.
@@ 2.2 Selective classification (SC)
A simple downstream task for estimates of predictive uncertainty is to reject (or detect) predictions that may incur a high cost (Xia & Bouganis, 2022b; Jaeger et al., 2023), using a binary rejection function,
@@
where $U(\boldsymbol{x})$ is a scalar measure of predictive uncertainty extracted from the prediction model and $\tau$ is a user-set operating threshold. Intuitively, we reject if the model is uncertain, preventing the system from serving failures, which may then be processed separately. We are only concerned with the relative rankings of $U$s rather than absolute values - we want to be more uncertain on failures. In the case where the prediction task is classification and we wish to reject potential misclassifications (✗), we can use a selective classifier (Chow, 1970; El-Yaniv & Wiener, 2010) $(f,g)$, which is simply the combination of a classifier $f$ (Eq. 5) and the aforementioned binary rejection function $g$ (Eq. 9). Fig. 2 contains an illustration. To evaluate a selective classifier, we use the 0/1 classification error,
@@
to define the selective risk (El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017) as
@@
$U(\boldsymbol{x})=-\text{MSP}(\boldsymbol{x})=-\pi_{\text{max}}=-\max\nolimits_{k}\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=-P(\hat{y}|\boldsymbol{x};\boldsymbol{\theta})~{},$ (12)
i.e. the (negative of the) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017) is the model’s estimate of the probability of its prediction $\hat{y}$. We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature (Xia & Bouganis, 2022b; 2023; Jaeger et al., 2023; Feng et al., 2023) compared to alternatives. MSP is a natural choice as $P(y|\boldsymbol{x};\boldsymbol{\theta})$ models $P_{\text{data}}(y|\boldsymbol{x})$, and $U=-\max_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ (or any scalar monotonic to it) provides the optimal risk if the true distribution $P_{\text{data}}(y|\boldsymbol{x})$ is known (Chow, 1970). We omit evaluations on other softmax-based $U$ from the main paper as we find them to behave similarly to MSP. In line with previous work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Zhu et al., 2024), we find that OOD detection (Yang et al., 2021) scores perform poorly at SC, and also omit them (see Sec. E.1 for discussion and additional results).
@@
Over/Underconfidence. Since we are only concerned about the relative ranking of $U$, we loosely refer to overconfidence as when a model’s estimate of uncertainty for a given sample is (relatively) low when $P_{\text{error}}$ high, as we want a model to be uncertain when it is likely to be wrong. Underconfidence is then the inverse. Both will result in worse SC: overconfidence leads to errors being accepted, underconfidence to correct predictions being rejected. Intuitively, over/underconfidence may arise from poorness of fit, where $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})$ fails to accurately model the true conditional distribution $\bar{\boldsymbol{\pi}}(\boldsymbol{x})$ (Fig. 4).
This is different to the definition used in model calibration (Guo et al., 2017), which, notably, is concerned with marginal properties (averaged over the input distribution) rather than per-sample properties, as well as absolute probabilities rather than relative rankings. It is possible to be highly over/underconfident in calibration but optimal for SC and vice versa (Zhu et al., 2024), e.g. applying the monotonic transformation $U=-\max_{k}[P_{\text{data}}(\omega_{k}|\boldsymbol{x})]^{1000}$. We note that model calibration is an independent task to SC (Jaeger et al., 2023), is well explored in the context of LS (Müller et al., 2019), and is not the focus of this work. LS may improve calibration by increasing the uncertainty of both correct ✓ and incorrect ✗ predictions, however, this will not necessarily help SC (Zhu et al., 2024).
3 The effect of label smoothing on selective classification
Zhu et al. (2022; 2024) observe empirically (as part of a broader investigation) that for a single value of $\alpha$ LS degrades SC for CNN-based image classification. In this section we aim to validate this behaviour on a wider range of large-scale tasks and architectures. We find that LS leads to consistent degradation in SC, with stronger LS leading to greater degradation. We suggest that it may partially explain the findings of recent large-scale empirical evaluations of open-source pre-trained models (Galil et al., 2023; Cattelan & Silva, 2024) where strong classifiers surprisingly underperform on SC.